# VLMQ: Efficient Post-Training Quantization for Large Vision-Language Models via Hessian Augmentation

### v1

@@ 1 Introduction
Large language models (LLMs) (Bubeck et al. 2023; Touvron et al. 2023a, b; DeepSeek-AI 2025; Jiang et al. 2023) have demonstrated exceptional advancements across diverse natural language processing tasks, leading to an increased emphasis on developing vision-language models (VLMs) (Wang et al. 2024b; Bai et al. 2025; Li et al. 2024b; Chen et al. 2024d) that process multi-modal inputs, including texts, images, and videos. Despite their impressive capabilities, the unprecedented scaling in the model size complicates their deployment in practical resource-limited contexts. Quantization methods have provided an effective solution by converting full-precision weights and activations (e.g., FP16/BF16) into reduced-precision formats (e.g., INT8/INT4), thereby markedly lowering memory footprint and computation complexity at the expense of minor performance degradation.
Figure 1: Baseline vs. VLMQ. The diagonal matrix in VLMQ represents the importance factors. The reported accuracy is from 2-bit Qwen2-VL-7B-Instruct (Wang et al. 2024b) quantized by GPTQ and VLMQ.
@@
Post-training quantization (PTQ) has emerged as a prevalent approach for deploying large-scale models, owing to its minimal computational overhead and the ability to bypass the costly fine-tuning or retraining process. Substantial research efforts have been directed toward designing advanced PTQ methods tailored LLMs. Such approaches typically aim to refine the distributions of weights and activations through various strategies, including equivalent transformations (Lin et al. 2023; Xiao et al. 2023), Hessian-based error compensation (Frantar and Alistarh 2022; Li et al. 2025b), and incoherence processing (Ashkboos et al. 2024; Hu et al. 2025). While significant progress has been achieved in applying PTQ to LLMs, their extension to VLMs has not yet been adequately investigated. Recent pioneering research efforts, such as MBQ (Li et al. 2025a), MQuant (Yu et al. 2025), and QSLAW (Xie et al. 2024), emphasize the modality imbalance challenge and propose dedicated importance-aware strategies to enhance the performance of quantized VLMs. However, these methods either require expensive parameter fine-tuning (Xie et al. 2024), specialized manipulation at inference time (Yu et al. 2025), or rely on a suboptimal grid search (Li et al. 2025a), failing to offer an efficient and effective solution for both calibration and inference.
Figure 2: PCA-based Hessian feature analysis with activations (4096 points) extracted from the pre-attention trace of the $20$-th transformer layer in Qwen2-VL-7B-Instruct. The left two subfigures depict the Hessian feature distributions constructed from text-only and mixed text-vision activations, respectively. The right three subfigures visualize the distributions with varying token-level importance factors. Light red/green and dark red/green points denote tokens classified as important or unimportant ones. Reported average accuracy is under INT3 quantization across eight vision-language benchmarks.
In this work, we identify a modality disproportion (i.e., limited text tokens vs. excessive and redundant vision tokens) in VLMs (Chen et al. 2024c; Xing et al. 2024; Alvar et al. 2025). We refer to redundancy as vision over-representation. Existing Hessian-based PTQ methods, such as GPTQ (Frantar et al. 2022), optimize the layer-wise reconstruction loss by treating all tokens equally. In these methods, the Hessian matrix that serves as a crucial guide for weight quantization is estimated by uniformly aggregating token contributions, regardless of their informativeness or importance. This token-agnostic formulation inherently biases the Hessian toward dominant yet redundant vision features, leading to sub-optimal weight updates. As a result, directly applying these methods to VLMs results in severe performance degradation.
To mitigate the side effects of vision over-representation, we propose VLMQ, an importance-aware calibration framework that optimizes a refined objective. As illustrated in Figure 1, VLMQ modifies the standard Hessian into an enhanced variant by assigning higher importance to salient tokens and lower importance to redundant ones. It yields a closed-form weight update while capturing token-level importance via the augmented Hessian. VLMQ is fully compatible with existing Hessian-based PTQ frameworks and readily inherits efficient tricks such as Cholesky decomposition. To derive effective importance factors, we first establish a theoretical connection between block-level loss perturbation and token-level quantization error. Based on this insight, we propose to perform a single lightweight block-wise backpropagation approach to derive layer-wise gradient-driven importance factors efficiently.
The core contributions are summarized as follows:
  • We identify a critical mismatch between the vision redundancy in VLMs and the token-agnostic objective of existing Hessian-based PTQ methods, which treat all tokens equally during quantization. We posit that this overlooked vision redundancy distorts the Hessian estimation, thereby explaining the poor performance of prior LLM PTQ methods when applied to VLMs.
  • We propose a PTQ scheme that leverages an enhanced Hessian derived from an importance-aware objective, which explicitly accounts for the redundancy among vision tokens and supports high degrees of parallelism.
  • We theoretically connect block-wise loss perturbation and first-order token-level activation errors, which motivates the use of a single lightweight block-wise backward pass to derive gradient-driven importance factors. This design achieves both efficiency and effectiveness in capturing the importance of tokens.
@@
  • Through extensive experimentation, we demonstrate the superior performance of our approach for VLMs, especially under ultra-low-bit quantization settings. For example, our method achieves state-of-the-art (SOTA) results and achieves up to 16.45% accuracy improvement.
2 Preliminary
@@
Notation. We follow most notations in GPTAQ (Li et al. 2025b). Specifically, general model parameters including $N$, $C_{i}$, $C_{o}$, and $\mathcal{V}$} denote the sequence length, in-channel number (hidden size), out-channel number (intermediate size), and the vocabulary set. We adopt the bold lowercase and uppercase letters to represent row-vectors and matrices, respectively. The letters with a hat symbol (e.g., $\hat{{\mathbf{x}}}$ or $\hat{{\mathbf{W}}}$) represent the quantized weights or activations noised by quantization errors. The linear operation is described as ${\mathbf{Y}}={\mathbf{W}}{\mathbf{X}}$ where ${\mathbf{W}}\in{\mathbb{R}}^{C_{o}\times C_{i}}$ and ${\mathbf{X}}\in{\mathbb{R}}^{C_{i}\times N}$ represent the weight and activation. The indexing rule is that ${\mathbf{W}}_{j,:}$ and ${\mathbf{W}}_{:,j}$ indicate the $j$-th row and $i$-th column of matrix $\mathbf{W}$} respectively. A negative index signifies the removal of a row in a matrix (e.g., ${\mathbf{X}}_{-1}\in{\mathbb{R}}^{(C_{i}-1)\times N}$).
@@
VLM architecture. The advanced development of VLMs (Wang et al. 2024b; Bai et al. 2025; Chen et al. 2024d; Li et al. 2024b) handles data in different modalities, including text, image, and video. The VLMs comprise three key components: a visual encoder, a vision-text projector, and a Language Model (LM) backbone. The mainstream decoder-only LM backbones output a distribution ${\mathbf{P}}=\left(p_{1},\;p_{2},\;...,\;p_{N-1}\right)\in{\mathbb{R}}^{N\times|{\mathcal{V}}|}$ and refer to it to generate predictions via diverse decoding strategies (Stern, Shazeer, and Uszkoreit 2018; Xia et al. 2022; Kim et al. 2022).
@@
We define input for each decoding layer in the above-mentioned LM backbone as activation ${\mathbf{X}}\in{\mathbb{R}}^{C_{i}\times N}$, where $C_{i}$ and $N$ are the model dimension (i.e., the input channel number) and sequence length, respectively. The activation transforms across $L$ decoding layers in LM backbones. The attention stream and feed-forward network stream within one decoding layer are defined as
@@
where $\texttt{MHSA}(\cdot)$ describes the multi-head self-attention operation with Q/K/V/O projections and $\texttt{FFN}(\cdot)$ represents the feed forward data flow with Up/Gate/Down projections. We omit the layer normalization involved in decoding layers for simplicity.
3 Related Work
PTQ for VLMs. The PTQ becomes challenging for VLMs due to dissimilar features and distributions between text and vision modality, and remains underexplored. MQuant (Yu et al. 2025) proposes modality-specific static quantization and attention-invariant flexible switching to address the modality gap and visual outliers. MBQ (Li et al. 2025a) observes that the error-sensitivity varies across text and vision tokens, and a corresponding biased objective is formulated to determine the optimal channel-wise equalization scales. QSLAW (Xie et al. 2024) determines the quantization step size via a learning-based method and proposes importance-aware warmup to alleviate the overfitting issue. Unlike these methods concentrating on modality difference, Q-VLM (Wang et al. 2024a) captures cross-layer dependency in VLMs and partitions blocks based on entropy.
OBQ, GPTQ, and GPTAQ. Hessian information is widely used as a calibration guide to compensate for quantization errors. OBQ calibrates the quantization of linear layers in models by iteratively 1) quantizing entries that introduce minimal loss, and 2) updating the remaining unquantized ones with the guidance of Hessian ${\mathbf{H}}={\mathbf{X}}{\mathbf{X}}^{\top}$. GPTQ (Frantar et al. 2022) extends this paradigm by formulating fixed-order entry quantization in a parallel manner, lazy block updating and Cholesky reformulation to achieve efficient calibration. GPTAQ (Li et al. 2025b) further extends GPTQ by optimizing an asymmetric objective:
[…]
where $q$ is the index of the quantized weight element. GPTAQ thus enables accurate and efficient quantization with better preservation of model functionality.
@@ 4 Motivation
Figure 3: Pipeline of computing importance factors. The “Forward” module illustrates the quantization dataflow across decoding layers, where a breakpoint is set at the output of each attention module to compute the local loss ${\mathcal{L}}_{\text{Block}}$ and trigger a localized backward pass. The “Backward” module details the internal operations within an attention block, where gradients of each linear projection output are cached to derive token-level importance factors. The “Process Grad” module summarizes how we formulate importance factors from raw gradients.
@@
Table 1: Statistics under controlled settings where a random subset of vision tokens are manually assigned as low-importance (LI) tokens. denotes the percentage of vision tokens whose importance values are set to 0.01, while the remaining tokens retain an importance value of 1.0.
Visual Over-Representation
The visual over-representation problem is a widely recognized issue in VLMs (Chen et al. 2024c; Xing et al. 2024; Alvar et al. 2025), and recent studies have explored vision token reduction to accelerate inference while preserving performance. However, existing Hessian-based PTQ methods, such as GPTQ (Frantar et al. 2022), fail to exploit this redundancy. They construct the Hessian by uniformly aggregating all token contributions that are derived from a standard optimization objective. As a result, these token-agnostic methods often struggle when applied to VLMs, where redundant vision tokens dominate the activation space. This limitation motivates the need for importance-aware quantization strategies that explicitly differentiate informative vision tokens from redundant ones.
Pilot Study
Pilot study on vision-role in quantization. To further investigate the role played by vision tokens in quantization performance, we conduct a pilot study that reveals how vision tokens influence Hessian estimation and model accuracy. We observe that 1) it is essential to incorporate vision tokens for effective quantization in VLMs, yet 2) an excessive number of redundant vision tokens may lead to degraded performance. Assigning redundant tokens with lower importance can mitigate the resulting performance degradation. To validate this insight, we benchmark INT3-quantized LLMs on DocVQA (Mathew, Karatzas, and Jawahar 2021) by randomly down-weighting a portion of vision tokens using a low importance factor. As shown in Table 1, model performance peaks when 50% of vision tokens are marked as low-importance (LI), which indicates that a balanced level of visual input is critical for optimal quantization.
Explanation of the pilot study. We attribute the observed performance fluctuation to the Hessian feature shift, as visualized in Figure 2. The leftmost subfigure shows that text-only tokens yield a compact and centralized distribution in the dominant component space. Introducing vision tokens, as in the second subfigure, diversifies the calibration distribution and helps mitigate the quantization–inference gap. However, due to the visual over-representation, excessive inclusion introduces the risk of Hessian bias toward redundant visual features.
We further color-code token points by their importance factors $\mathbf{G}$} (defined later) in the right three subfigures. The evident imbalance in importance density across modalities risks skewing the Hessian toward redundant visual margins, rather than concentrating on the more informative central regions. These findings highlight the necessity of fine-grained importance-aware quantization to counteract redundancy-induced bias.
Overall, our pilot study highlights the necessity of incorporating visual input, whereas overabundance impairs quantization quality. The performance peak observed with a moderate down-weighting of vision tokens validates the importance of balanced visual input. These insights motivate the need for a fine-grained, importance-aware quantization strategy that selectively emphasizes salient vision tokens while mitigating the side effects of redundant ones.

### v2

# VLMQ: Token Saliency-Driven Post-Training Quantization for Vision-language Models

@@ Abstract
footnotetext: $\dagger$} Work done during an internship at TeleAI.footnotetext: $\ddagger$} Corresponding authors.
@@ 1 Introduction
Large language models (LLMs) [bubeck2023gpt4, touvron2023llama, touvron2023llama2, deepseek2025deepseek-r1, jiang2023mistral] have demonstrated exceptional advancements across diverse natural language processing tasks, leading to an increased emphasis on developing vision-language models (VLMs) [wang2024qwen2vl, bai2025qwen25vl, li2024llava-ov, chen2024internvl2, xiaomi2025mimovl, zhu2025internvl3] that process multi-modal inputs, including texts, images, and videos. Despite their impressive capabilities, the unprecedented scaling in the model size complicates their deployment in practical resource-limited contexts [an2026aiflow, shao2025aiflow, yuan2025task].
@@
In light of these problems, quantization has provided an effective solution by converting full-precision weights and activations (e.g. FP16/BF16) into reduced-precision formats (e.g. INT8/INT4), thereby markedly lowering memory footprint and computation complexity. To be more specific, post-training quantization (PTQ) has emerged as a prevalent approach for deploying large-scale models, owing to its minimal computational overhead and the ability to bypass the costly fine-tuning or retraining process. Substantial research efforts have been directed toward designing advanced PTQ methods tailored to LLMs. Such studies typically aim to refine the distributions of weights and activations through various strategies, including equivalent transformations [lin2023awq, xiao2023smoothquant], Hessian-based error compensation [frantar2022optimal-brain-compression, li2025gptqv2], and incoherence processing [ashkboos2024quarot, hu2025ostquant]. While significant progress has been achieved in applying PTQ to LLMs, their extension to VLMs has not yet been adequately investigated. Recent pioneering research efforts, such as MBQ [li2025mbq], MQuant [yu2025mquant], and QSLAW [xie2024qslaw], emphasize the modality imbalance challenge and propose dedicated importance-aware strategies to enhance the performance of quantized VLMs. However, these methods either require expensive parameter fine-tuning [xie2024qslaw], specialized manipulation at inference time [yu2025mquant], or rely on a suboptimal grid search [li2025mbq], failing to offer an efficient and effective solution for both calibration and inference.
Figure 1: Baseline vs. VLMQ. The diagonal matrix in VLMQ represents the importance factors. The reported accuracy is from 2-bit Qwen2-VL-7B-Instruct [wang2024qwen2vl] quantized by GPTQ and VLMQ.
In this work, we highlight two intrinsic properties of VLMs that critically affect quantization performance. In particular, VLM activations reveal 1) a pronounced visual over-representation (i.e. limited text tokens vs. excessive and redundant vision tokens) as well as 2) a modality gap. However, existing PTQ approaches for LLMs minimize the layer-wise reconstruction loss while treating all tokens uniformly [lin2023awq, frantar2022gptq, li2025gptqv2], without accounting for token-level informativeness or importance. Such a token-agnostic design inevitably biases the quantized model toward dominant but redundant visual features, thereby yielding significant performance drops. Consequently, directly transferring these methods to VLMs is suboptimal.
To alleviate the adverse effects of visual over-representation and modality gap, we propose VLMQ, an importance-aware calibration framework specifically designed for VLMs. We introduce an importance factor $\mathbf{G}$} to capture token-wise variations in informativeness. To obtain effective importance factors, we first formulate a theoretical and empirical connection between loss perturbation and token-level quantization error. Guided by this Theorem, we further adopt a lightweight block-wise backpropagation strategy that efficiently derives gradient-driven, layer-specific importance factors. Finally, as shown in Figure 1, VLMQ refines the conventional optimization objective into an importance-aware formulation, assigning higher weight to salient tokens while down-weighting redundant ones. The core contributions are summarized as follows:
  • $\triangleright$
    We uncover a fundamental mismatch between the vision redundancy inherent in VLMs and the token-agnostic objectives of existing mainstream PTQ approaches, which treat all tokens uniformly during quantization. We verify that this neglected redundancy disrupts layer reconstruction, thereby accounting for the degraded performance observed when directly applying LLM PTQ methods to VLMs.
  • $\triangleright$
    We propose a gradient-driven importance factor $\mathbf{G}$} to suppress less informative tokens. Its effectiveness is theoretically and experimentally supported through the established link between loss perturbation and first-order token-level activation errors, while its efficiency is ensured via a lightweight block-wise backpropagation scheme.
  • $\triangleright$
@@
    Through extensive experiments, we validate the superior performance of our framework on VLMs, particularly under ultra-low-bit quantization regimes. Notably, our approach achieves state-of-the-art (SOTA) results and yields accuracy improvements of up to 16.45%.
2 Background
2.1 Preliminary
2.1.1 Notation.
@@ 2.1 Preliminary
NN, $C_{i}$, $C_{o}$, and $\mathcal{V}$} denote the sequence length, in-channel number (hidden size), out-channel number (intermediate size), and the vocabulary set. We adopt the bold lowercase and uppercase letters to represent row vectors and matrices, respectively. The letters with a hat symbol (e.g. $\hat{{\mathbf{x}}}$ or $\hat{{\mathbf{W}}}$) represent the quantized weights or activations noised by quantization errors. The linear operation is described as ${\mathbf{Y}}={\mathbf{W}}{\mathbf{X}}$ where ${\mathbf{W}}\in{\mathbb{R}}^{C_{o}\times C_{i}}$ and ${\mathbf{X}}\in{\mathbb{R}}^{C_{i}\times N}$ represent the weight and activation. The indexing rule is that ${\mathbf{W}}_{j,:}$ and ${\mathbf{W}}_{:,j}$ indicate the $j$-th row and $i$-th column of matrix $\mathbf{W}$} respectively. A negative index signifies the removal of a row in a matrix (e.g. ${\mathbf{X}}_{-1}\in{\mathbb{R}}^{(C_{i}-1)\times N}$).
2.1.2 VLM architecture.
@@
The advanced development of VLMs [wang2024qwen2vl, bai2025qwen25vl, chen2024internvl2, li2024llava-ov] handles data in different modalities, including text, image, and video. The VLMs comprise three key components: a visual encoder, a vision-text projector, and a language model (LM) backbone. The mainstream decoder-only LM backbones output a distribution $\mathbf{Prob}=\left(p_{1},\;p_{2},\;...,\;p_{N-1}\right)\in{\mathbb{R}}^{N\times|{\mathcal{V}}|}$ and refer to it to generate predictions via diverse decoding strategies [stern2018blockwise, xia2022speculative, kim2022critic].
@@
We define input for each decoding layer in the above-mentioned LM backbone as activation ${\mathbf{X}}\in{\mathbb{R}}^{C_{i}\times N}$. The activation transforms across $L$ decoding layers in LM backbones. The attention stream and feed-forward network stream within one decoding layer are defined as
@@
where $\texttt{MHSA}(\cdot)$ describes the multi-head self-attention operation with Q/K/V/O projections and $\texttt{FFN}(\cdot)$ represents the feed-forward data flow with Up/Gate/Down projections. We omit the layer normalization involved in decoding layers for simplicity.
2.2 Related Work
2.2.1 Advanced VLMs.
Building upon the rapid advancements in LLMs, VLMs have demonstrated exceptional capabilities in visual understanding [li2024llava-ov, wang2024qwen2vl, bai2025qwen25vl, xiaomi2025mimovl, chen2024internvl2, zhu2025internvl3]. For instance, Qwen2-VL [wang2024qwen2vl] enhances multimodal representations through techniques such as M-RoPE and naive dynamic resolution, facilitating unified text-image-video comprehension with flexible visual tokenization. Qwen2.5-VL [bai2025qwen25vl] further refines VLM capabilities by enabling omnidocument parsing, supporting ultra-long video understanding, and enhancing agent functionalities across both desktop and mobile platforms, thereby exhibiting advanced vision-language intelligence. InternVL3 [zhu2025internvl3] contributes to the evolution of multimodal large language models (MLLMs) by introducing variable visual position encoding for handling extended contexts. Leveraging these innovations, it excels in multimodal perception, reasoning, and textual performance.
2.2.2 PTQ for large models.
PTQ has become a widely adopted technique for compressing [li2024tesseraq, wnag2024ptsbenchcomprehensiveposttrainingsparsity, huang2025qvgenpushinglimitquantized, gong2024llmcbenchmarkinglargelanguage] and accelerating large models [lv2024ptq4sam, chen2024db-llm, tian2024qvd, huang2025harmonicaharmonizingtraininginference]. In the context of LLMs, approaches such as AWQ [lin2023awq] and SmoothQuant [xiao2023smoothquant] address activation outliers through smooth-based transformations. Hessian-informed methods [frantar2022optimal-brain-compression, frantar2022gptq, li2025gptqv2] provide closed-form, layer-wise quantization schemes. More recently, rotation-based methods [ashkboos2024quarot] have improved quantization robustness by reshaping parameter distributions, thereby improving low-bit quantization performance. However, the PTQ becomes challenging for VLMs due to dissimilar features and distributions between text and vision modality, and remains underexplored. MQuant [yu2025mquant] proposes modality-specific static quantization and attention-invariant flexible switching to address the modality gap and visual outliers. MBQ [li2025mbq] observes error-sensitivity variance across different modalities. Built on [lin2023awq], an improved scale searching strategy is proposed that takes modality variance into consideration. QSLAW [xie2024qslaw] determines the quantization step size via a learning-based method and proposes modality-aware warmup to alleviate the overfitting issue. Unlike these methods concentrating on modality difference, Q-VLM [wang2024q-vlm] captures cross-layer dependency in VLMs and partitions blocks based on entropy.
@@ 3 Motivation
Figure 2: PCA-based [shlens2014pca] activation feature analysis with activations (4096 points) extracted from the pre-attention breakpoint of the $20$-th transformer layer in Qwen2-VL-7B-Instruct. The left two subfigures depict the activation feature distributions constructed from text-only and mixed text-vision activations, respectively. The right three subfigures visualize the distributions with varying token-wise importance factors. Light red/green and dark red/green points denote tokens classified as important or unimportant ones. Reported average accuracy is under INT3 quantization across eight vision-language benchmarks.
VLMs process inputs from both text and vision modalities, which exhibit significantly different statistical properties and distributional characteristics. We identify two key observations (i.e. visual over-representation and modality gap) that together have a substantial impact on the performance of quantized models. In this section, we first introduce these two observations and discuss the insights they provide. Then, we present a pilot study to empirically validate these findings.
3.1 Modality Discrepancy
3.1.1 Observation 1: visual over-representation.
We have found that VLM inputs contain an excessive number of redundant vision tokens, which dominate the activation space (see Figure 2). We term this phenomenon as visual over-representation. Current PTQ approaches originally designed for LLMs overlook this redundancy. They minimize the layer-wise mean square error (MSE) by treating all tokens equally, following a standard optimization formulation [lin2023awq, frantar2022gptq, li2025gptqv2, shao2023omniquant].
3.1.2 Observation 2: modality gap.
We also uncover a modality gap in VLMs, which denotes the distinct separation between text tokens and vision tokens in the latent feature space, as shown in Figure 2. Such misalignment can bias the calibration process, especially for existing studies [lin2023awq, frantar2022gptq], which disproportionately favor excessive and redundant vision tokens over informative ones during quantization. As a result, existing token-agnostic methods often incur non-negligible accuracy drops in VLMs (see Section 5).
3.2 Pilot Study
3.2.1 Insight: down-weighting vision tokens improves quantization.
These observations motivate the need for importance-aware quantization strategies that explicitly distinguish salient vision tokens from redundant ones. We hypothesize that assigning lower weights to a subset of vision tokens balances the distribution density across modalities, thereby enhancing quantization performance. To validate this hypothesis, we conduct the following pilot study.
Table 1: Statistics under controlled settings where a random subset of vision tokens are manually assigned as low-importance (LI) tokens. denotes LI tokens are down-weighted by a factor 0.01.
@@ 3.2 Pilot Study
3.2.2 Pilot study on vision-role in quantization.
To examine the impact of vision tokens on quantization, we conduct a pilot study assessing how they affect model performance. We benchmark INT3-quantized VLMs on DocVQA [mathew2021docvqa] by randomly down-weighting a subset of vision tokens with a low importance factor. As shown in Table 1, increasing the LI ratio within a certain range alleviates performance degradation. Notably, performance reaches its peak when 50% of vision tokens are marked as low-importance (LI), suggesting that maintaining a balanced level of visual input is crucial for effective quantization. To better explain this phenomenon, in Figure 2, we further color-code token points by their importance factors $\mathbf{G}$} (defined later) in the right three subfigures. The evident imbalance in importance density across modalities risks skewing the distribution toward redundant visual margins, rather than concentrating on the more informative central regions. These findings highlight the necessity of fine-grained importance-aware quantization to counteract redundancy-induced bias.
Overall, the observed performance peak under moderate down-weighting of vision tokens underscores the importance of maintaining a balanced level of inputs. These findings further motivate the development of a fine-grained and importance-aware quantization strategy that selectively prioritizes salient tokens while suppressing the influence of redundant ones. In the following section, we introduce VLMQ, a dedicated quantization framework for VLMs that leverages token-wise importance to counteract redundancy-induced quantization bias.